Find the solution of the following differential equation using Laplace transforms

y''+4y=e^{t}, y(0)=0, y'(0)=0
Answer #1

Given y+4y= et 90-0, y(0)=0 Apply Laplace transformes [py(D) – Py(9–y@] +49 (P= p-l py ® +49 (17) P| - y(D) - y (P= 5-D (P